Solve by elimination method:
3x-2y=20
and
2x+3y=-4

Answer by stanbon(75887)

Solve by elimination method:
3x-2y=20
and
2x+3y=-4
-----------
Multiply thru the 1st by 2.
Multiply thru the 2nd by 3.
-------------------------
6x - 4y = 40
6x + 9y = -12
-----------
Subtract and solve for "y":
13y = -52
y = -4
------
Solve for "x":
----------
2x+3y = -4
2x -12 = -4
2x = 8
x = 4
----
Solution (4,-4)
=====================
Cheers,
Stan H.